Find all real and imaganiary solutions to each equation
b^4 + 13b^2 + 36 = 0
Thanks for the help !

Answer by stanbon(75887)

Find all real and imaginary solutions to each equation
b^4 + 13b^2 + 36 = 0
---
Factor to get:
(b^2+9)(b^2+4) = 0
Factor to get:
(b+3i)(b-3i)(b+2i)(b-2i) = 0
a = -3i or 3i or -2i or 2i
=======================================
Cheers,
Stan H.